Two angles are complementary. The measure of one angle is one-fifth of the measure its complement. Find the measure of the smaller angle.

Please answer this Thank you.Have a nice day. :]

Answer by fcabanski(1391)

Check my profile and contact me for online tutoring.


Complimentary angles add up to 90 degrees.


Call the first angle x, and the second (1/5)x. Their sum is 90.


x + (1/5)x = 90 (Add like terms).


(6/5)x=90 (divide both sides by 6, multiple both sides by 5).


x = 90*(5/6) = 75


X is 75, the other angle is (1/5)*75 = 15 degrees.